OpenStudy (anonymous):

Find the solutions of the inequality. | x - 5 | ≥ 3

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x \ge 8\ or\ x \le 2\]

OpenStudy (anonymous):

can you show me how to sovle this equation

OpenStudy (anonymous):

thank you for the answer

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

You should see it as two cases: "Case 1" x-5 <= 3 (when x> = 0) "Case 2": - (x-5) <= 3 (when x <0) Then count the cases separately... Write back so I explain more if you want ...